OpenStudy (anonymous):

A woman drives 10 miles, accelerating uniformly from rest to 70 mph. How long does it take for her to reach 70 mph?

OpenStudy (anonymous):

v = u + at v = her final velocity u = her initial velocity a = acceleration t = time take (which you are finding)

OpenStudy (anonymous):

what's acceleration in this case?

OpenStudy (anonymous):

oh sorry. u need a different equation

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

first find her acceration \[v ^{2} = u ^{2} + 2as\] then use v = u + at

OpenStudy (anonymous):

for kinematics, you need 3 equations and u can solve everything

OpenStudy (anonymous):

ill post them for you. copy them somewhere

OpenStudy (anonymous):

\[v = u + at\] \[v ^{2} = u ^{2} + 2as\] \[s = ut + \frac{ 1 }{ 2 }at ^{2}\]

OpenStudy (anonymous):

ok what's s?

OpenStudy (anonymous):

s = displacement remember these are vectors so direction matters as that decides whether the value is positive or negative